"The weights of 1,000 children, in average, is 62kg with a standard deviation of 16kg. Suppose the weights are normally distributed, how many children weigh less than 51kg?"




1
Expert's answer
2021-02-24T07:00:17-0500

Given : mean= 62

sd = 16


formula Z= "\frac{x-mean}{sd}"

solution:

For x= 51, Z= "\frac{51-62}{16}" = -0.6875


Using Z table , Pr(Z≤−0.6875)=0.2459

Hence number of children = 1000*0.2459= 245.9 approximately 246 children